# ByteGen: A Tokenizer-Free Generative Model for Orderbook Events in Byte Space

### v1

@@ 3 Methodology
Our methodology is based on a novel hierarchical architecture designed to learn directly from raw, untokenized data streams (hwang2025dynamic, ). The core idea is to build a model that mimics a multi-level reasoning process, moving from fine-grained details to high-level abstractions, all within a single, end-to-end system.
@@ 3.1 Data Processing to Transform Raw Market Data to Byte Data
[…]
@@
[…]
@@
$\text{iids}=\text{NJT}([s[:-1]\text{ for }s\text{ in samples}]),\quad\text{lbls}=\text{NJT}([s[1:]\text{ for }s\text{ in samples}])$ (2)
@@ 3.2 Hierarchical Neural Transformer (H-Net) with Dynamic Chunking
where $F$ is the fraction of selected boundaries, $G$ is the average boundary probability, and $N$ is the target compression ratio. We use compression ratios of [1,4,16][1,4,16][ 1 , 4 , 16 ] across stages.
@@ 3.3 Mamba Architecture
[…]
@@ 3.4 Transformer Architecture
[…]
@@ 5.1 Future Work
Finally, the current 32-byte fixed format, while efficient, limits applicability to exchanges with different event encodings. Future work could explore variable-length event formats through architectural modifications, potentially using length-prefixed encodings or learnable event boundaries. This would enable ByteGen to process data from multiple exchanges without format conversion, moving toward a universal market data model. This study can be used in application such as (li2025flowhft, ; li2025flowoe, ).

### v2

@@ 3 Methodology
Our methodology is based on a novel hierarchical architecture designed to learn directly from raw, untokenized data streams (hwang2025, ). The core idea is to build a model that mimics a multi-level reasoning process, moving from fine-grained details to high-level abstractions, all within a single, end-to-end system.
@@ 3.1 Data Processing to Transform Raw Market Data to Byte Data
$\text{Event}=\text{Pack}(\text{ev\_packed},\text{exch\_ts},\text{price},\text{quantity})$ (1)
@@
This packed representation reduces the original 64-byte format (8 fields) to 32 bytes while preserving all essential information. The bitwise packing scheme $\text{ev\_packed}=(\text{order\_id}\ll 32)\mid\text{ev}$ enables efficient storage and processing.
@@
[…]
@@ 3.2 Hierarchical Neural Transformer (H-Net) with Dynamic Chunking
where $F$ is the fraction of selected boundaries, $G$ is the average boundary probability, and $N$ is the target compression ratio. We use compression ratios of $[1,4,16]$ across stages.
@@ 3.3 Mamba Architecture
In H-Net, it uses the Mamba-2. The process can be broken down into the following steps. The first is input expansion. The process begins with a linear projection that expands the dimensionality of the input sequence, $u$. Typically, the dimension is doubled. This expanded representation is then split into two independent streams, $x$ and $z$, which will follow parallel paths through the block. This initial expansion provides the model with a richer, higher-dimensional space to perform its computations. The stream $x$ undergoes a series of transformations before being processed by the selective scan. First, $x$ is passed through a $1D$ causal convolution. The purpose of this step is to capture immediate local context. It allows the model to gather information from a small, fixed-size window of recent inputs before feeding them into the state model. This helps the SSM by pre-processing the input to account for short-range patterns. The output of the convolution is then fed into linear layers that generate the input-dependent parameters for the SSM: the matrices $B$ and $C$, and the timestep $\Delta$. This is the core of the selection mechanism, where the model learns to dynamically shape its own parameters based on the local context. Using the generated $B_{k}$, $C_{k}$, and $\Delta_{k}$ , along with the fixed state transition matrix $A$, the model updates its hidden state $h_{k}$ and computes the output for this path using the highly efficient parallel scan algorithm. The second stream, $z$, follows a much simpler path. It is passed through a non-linear activation function, typically SiLU (Sigmoid-Weighted Linear Unit), also known as Swish. This path acts as a gating mechanism, similar to the gates in an LSTM or GRU. Its purpose is to learn which information from the main SSM path is actually relevant and should be passed on. The output of the SSM path is then modulated by the gating path via element-wise multiplication (Hadamard product). The output of the SiLU activation on stream z effectively acts as a filter, scaling the output of the SSM. If the gate’s output is close to zero for certain dimensions, the corresponding information from the SSM is suppressed. If it is large, the information is passed through. Finally, this modulated result is passed through a linear layer to project it back down to the original input dimension.
@@ 3.4 Transformer Architecture
This computation can be broken down into four steps: The first is to calculate compatibility scores. The term $QK^{T}$ computes the dot product between every query vector and every key vector. This results in an attention matrix of size $L\times L$, where each element $(i,j)$ represents the compatibility or relevance of token $j$ to token $i$. The second step is for scaling. These scores are then scaled by dividing by $\sqrt{d_{k}}$, the square root of the dimension of the key vectors. This scaling factor is crucial for stabilizing the training process, as it prevents the dot products from becoming too large, which could push the softmax function into regions with extremely small gradients. The third step is to do normalization. A softmax function is applied row-wise to the scaled attention scores. This converts the raw compatibility scores into a probability distribution, ensuring that for each token, the attention weights assigned to all other tokens in the sequence sum to 1. Finally, the resulting attention weights are multiplied by the Value matrix $V$. This step produces the final output for each token, which is a weighted sum of all Value vectors in the sequence. In essence, the representation for each token becomes a rich, context-aware embedding that has selectively aggregated information from the entire sequence based on the learned attention patterns.
@@ 5.1 Future Work
Finally, this research shows promise for applications in several fields, as illustrated by recent studies (li2025flowhft, ; li2025flowoe, ). These works generate action segments based on sequences of past observations by mapping obs[t-1,t] to action[t-1,t,t+1,…]. Moving forward, research might employ ByteGen to forecast future observations $\mathrm{obs\_pred}[t+1]$, thus enhancing the mapping from obs[t-1,t,t+1] to action[t-1,t,t+1,…]. By incorporating predicted future states, this approach could improve action generation performance by offering the model a more comprehensive temporal framework.